# Evaluating Open Language Models Across Task Types, Application Domains, and Reasoning Types: An In-Depth Experimental Analysis

### v1

@@ Appendix E Prompt Analyses
E.1 Best Prompt Style at LM Level
In Table 1, we reported the best prompt style at the LM-level, abstracting all analyses at aspect-level. To support that, we report BERTScore recall values for all prompt styles used in this work at Language Model level without going into the aspects in Table 12. These are scores on the entire experimental dataset.
@@ E.1 Best Prompt Style at LM Level
From the table, we see that the differences with increasing examples are less prominent than observed at aspect-level analyses of prompt style in Section 3.5 and Appendix E.2. Therefore, an aspect-level analysis of performance is important to determine the prompt style for a given use-case.
E.2 Variation of Performance with Different Prompt Styles for all Language Models
This is a continuation from Section 3.5 where we analyzed how performance of Mistral-7B-I varied for different task types, domains and reasoning types with the 8 different prompt styles that we use. In this section, we will provide similar visualizations for all other models. Using these graphs, not only one can determine the best prompt style for that particular task type, domain, or reasoning type, but if using that is not possible, the performance trade-off of using any other prompt style can also be analyzed. The visualizations are provided in Figures 8 – 17. From these, it is clearly visible that for each LM, the variation in performance is different for each entity of task type, application domain and reasoning type. Therefore, the prompt style should be carefully selected by examining the trend.
E.3 Praphrasing Definitions
@@
E.4 Adversarial Definitions
This experiment aims to identify how robust are these LMs when they are asked to complete a task instance with a task definition that has subtle differences that can confuse the model, or are meant to elicit a response that is not desired. These subtle differences can both be intentional and non-intentional.
To perform this, similar to Appendix E.3, we generate adversarial task definitions for all the task definitions available in the dataset using gpt-3.5-turbo (Brown et al., 2020; OpenAI, 2023) using a pre-determined prompt which instructs it to generate the adversarial definition. The prompt is given below, and some examples are given in Table 15.
@@ E.4 Adversarial Definitions
Then, we use the prompt style with definition and 0 examples, but replace the definition with the adversarial definition of the task. At last, we calculate the BERTScore recall values for adversarial versus actual task definition. The reason to choose 0 examples was to avoid the scenario of the model recovering by learning from in-context examples.
@@
From the table, we see that most models are robust to adversarial change in task definition, with 7/10 models suffering less than 3% of decrease in performance. For the pre-trained model, Falcon-2-11B is most robust with only 0.613% decrease, and Gemma-2B ranks second with 0.88% decrease. Mistral-7B suffers a severe degradation in performance. For IT models, Gemma-2B-I is still one of the best, suffering only 1.2% decrease in BERTScore recall values only, but is outperformed by Llama-3-8B-I. Mistral-7B-I, the best performing IT model on true definitions is also not very sensitive to this change. Ph-3-mini-128k-I suffers the most amongst all models. We have seen sensitivity to be a general trend in this model with all varying parameters.
@@
(a) Task Types; with definition
(b) Domains; with definition
(c) Reasonings; with definition
(d) Task Types; without definition
(e) Domains; without definition
(f) Reasonings; without definition
Figure 9: Mean BERTScore recall variation for Mistral-7B across various task types, domains, and reasoning types by varying number of in-context examples, segmented by with and without using task definitions (Columns: different aspects, row 1: with task definition, row 2: without task definitions).
@@
Figure 14: Mean BERTScore recall variation for Phi-3-mini-128k-I across various task types, domains, and reasoning types by varying number of in-context examples, segmented by with and without using task definitions (Columns: different aspects, row 1: with task definition, row 2: without task definitions).
@@
Appendix F Implementation Details
@@
As discussed before, we are also sharing an anonymous GitHub repository  [^6] https://github.com/neelabhsinha/lm-application-eval-kit of our implementation as a utility which will allow evaluating any LM using this dataset, generate these visualizations, with other evaluation metrics also that are discussed in Table 1.
@@
Phi-3-mini-128k-I Model Card MIT License

### v2

@@ Appendix D Prompt Analyses
D.1 Best Prompt Style at LM Level
We report BERTScore recall values for all prompt styles used in this work at Language Model level without going into the aspects in Table 8. These are scores on the entire experimental dataset.
@@ D.1 Best Prompt Style at LM Level
From the table, we see that the differences with increasing examples are less prominent as compared to the aspect-level analyses of prompt style in Section 3.6 and Appendix D.2. This highlights the importance of conducting the prompt style analysis at aspect level. It is important to determine the prompt style that serves the best for a given use-case.
D.2 Variation of Performance with Different Prompt Styles for all Language Models
This is a continuation from Section 3.6 where we analyzed how performance of Mistral-7B-I varied for different task types, domains and reasoning types with the 8 different prompt styles that we use. In this section, we will provide similar visualizations for all other models. Using these graphs, one can determine the best prompt style for that particular task type, domain, or reasoning type. Additionally, the performance trade-off of using any other prompt style can also be analyzed. The visualizations are provided in Figures 8 – 17. From these, it is clear that for each LM, the variation in performance is different for each entity of task type, application domain and reasoning type. Therefore, the prompt style should be carefully selected by examining the trend.
D.3 Paraphrasing Definitions
@@
D.4 Adversarial Definitions
This experiment aims to identify how robust the LMs are when they are asked to complete a task instance with a task definition that has subtle differences capable confuse it, or are provided to elicit a response that is not desired. These subtle differences can both be intentional and non-intentional.
To perform this, similar to Appendix D.3, we generate adversarial task definitions for all the task definitions available in the dataset using gpt-3.5-turbo (Brown et al., 2020; OpenAI, 2023) using a pre-determined prompt which instructs it to generate the adversarial definition. The prompt is given below, and some examples are given in Table 11.
@@ D.4 Adversarial Definitions
Then, we use the prompt style with definition and 0 examples, but replace the definition with the adversarial definition of the task. At last, we calculate the BERTScore recall values for adversarial versus actual task definition, and report the results in Table 12. The reason to choose 0 examples was to avoid the scenario of the model recovering by learning from in-context examples.
@@
From the table, we see that most models are robust to adversarial change in task definition, with 7/10 models suffering less than 3% of decrease in performance. For the pre-trained model, Falcon-2-11B is most robust with only 0.613% decrease, and Gemma-2B ranks second with 0.88% decrease. For IT models, Gemma-2B-I is still one of the best, suffering only 1.2% decrease in BERTScore recall values only, but is outperformed by Llama-3-8B-I. Mistral-7B-I, the best performing IT model on true definitions is also not very sensitive to this change. Ph-3-mini-128k-I suffers the most amongst all models. We have seen sensitivity to be a general trend in this model with all varying parameters.
(a) Task Types; with definition
(b) Domains; with definition
(c) Reasonings; with definition
(d) Task Types; without definition
(e) Domains; without definition
(f) Reasonings; without definition
Figure 8: Mean BERTScore recall variation for SmolLM-1.7B across various task types, domains, and reasoning types by varying number of in-context examples, segmented by with and without using task definitions (Columns: different aspects, row 1: with task definition, row 2: without task definitions).
@@
Figure 13: Mean BERTScore recall variation for SmolLM-1.7B-I across various task types, domains, and reasoning types by varying number of in-context examples, segmented by with and without using task definitions (Columns: different aspects, row 1: with task definition, row 2: without task definitions).
@@
Appendix E Comparison of Decoding Techniques
Model Name Greedy top-k top-p
$k=10$ $p=0.9$
SmolLM-1.7B 83.71 83.72 83.71
Gemma-2B 86.41 86.16 86.38
Gemma-7B 81.06 81.02 81.02
Meta-Llama-3-8B 82.73 83.12 83.05
Falcon-2-11B 86.18 86.12 86.12
SmolLM-1.7B 86.66 86.64 86.67
Gemma-2B-I 87.96 88.01 87.91
Mistral-7B-I-v0.3 88.29 88.29 88.29
Gemma-7B-I 85.14 85.15 85.15
Meta-Llama-3-8B-I 84.11 84.12 84.18
Appendix F Qualitative Analyses of Generated Outputs
F.1 Qualitative Examples for Mistral-7B-I
Task Type Domain R. Type Instance Candidate Reference
F.2 Explaining Incorrect Responses of Pre-trained models
In Section 3.5 and Appendix B, we observed that even the best pre-trained models are not able to match the performance of IT models on SOTA models. While there was a theoretical reasoning, in this appendix, we will take some qualitative observation of outputs and focus on explainability of errors for the two best performing pre-trained models - Gemma-2B and Falcon-2-11B using the best prompt style.
Appendix G Implementation Details
@@
As discussed before, we are also sharing a GitHub repository of our implementation (link available on page 1 footnote) as a utility which will allow evaluating any LM using this dataset and generating these visualizations. It also supports doing this using other evaluation metrics discussed in Table 7 if required.
@@
Gemma-2-2B-I Model Card Apache 2.0 License